A factory foreman earns 11% more than their assistant. Together they make $119,426 per year. What is the assistant's annual salary (in dollars)?

Found 2 solutions by Theo, mananth:
Answer by Theo(13342)

let x equal the assistant's salary.
then 1.11 * x = the foreman's salary.
you get x + 1.11x = 119,426
combine like terms to get 2.11x = 119,426
solve for x to get x = 56,600.
that would be the salary of the assistance.
the forman makes 1.11 * that = 62,826
that should be your solution.
their combines salaries are 56,600 + 62,826 = 119,426.
this confirms their salaries are correct.

Answer by mananth(16949)

A factory foreman earns 11% more than their assistant. Together they make $119,426 per year. What is the assistant's annual salary (in dollars)?
Let assistant get $ x
the foreman will get 1.11x
Total salary of both = 119426
1.11x +x = 119426
2.11x = 119426
x = 119426/2.11
x = 56600
the assistant's annual salary is $56,600
Check
56600 +1.11*56600 = 119426